# Bias Amplification: Language Models as Increasingly Biased Media

### v1

@@ 4.1 Experiment Dataset Preparation
We randomly selected 1,518 articles from the Webis-Bias-Flipper-18 dataset (Chen et al., 2018), which contains political articles from a range of U.S. media outlets published between 2012 and 2018, along with bias ratings assigned at the time for each media source. These ratings were provided by Allsides and were determined through a multi-stage process involving both bipartisan experts and members of the general public (AllSides, 2024a). Based on the bias ratings, the random selection was stratified to ensure that the 1,518 articles were evenly divided into three groups of 506 articles each, representing left-leaning, right-leaning, and center-leaning media.
4.2 Fine-tuning and Synthetic Data Generation
@@
Inspired by Shumailov et al. (2024); Dohmatob et al. (2024b), each training cycle begins with GPT-2 fine-tuned on the dataset of human-written political articles. The fine-tuned GPT-2 is then used to generate the first set of synthetic data, which is subsequently employed to further fine-tune the model, producing model generation 1. This iterative process continues until generation 10 is reached. For details on the fine-tuning setup, refer to Appendix B.
[…]
@@ 4.3 Political Bias Metric
We develop a metric model to classify the political leaning of synthetic outputs generated by large language models. The classifier assigns labels corresponding to left, center, and right-leaning, trained on data from AllSides (AllSides, 2024a), a source that assigns bias ratings to each media outlet. The methodology for the classifier is as follows:
Dataset and Preprocessing. We use a dataset consisting of political articles with labeled bias categories for the source media: left, center, and right. We concatenate the article titles with their body text to form the input. To address class imbalance, we resample the center-leaning articles, balancing the dataset with respect to each category. The dataset is then split into training (70%), validation (15%), and test (15%) subsets, stratified by bias label. Moreover, we conduct a human review of the test subsets to eliminate any identifiable information about media sources or writers.
Model Architecture. We experiment with multiple transformer-based models, including BERT (Devlin et al., 2018), RoBERTa (Liu et al., 2019), and other state-of-the-art architectures, to select the best-performing model based on macro F1 score. Each model is fine-tuned using the HuggingFace ‘Trainer‘ class, with a learning rate of $2\times 10^{-5}$, a batch size of 16, and 5 training epochs. We employ a cross-entropy loss function for multi-class classification.
Training Procedure. Tokenization is performed using each model’s respective tokenizer with a maximum sequence length of 512 tokens. To mitigate overfitting, weight decay of 0.01 is applied during training. The model checkpoints are saved after each epoch, and the best model is selected based on macro F1 score evaluated on the validation set. We use a weighted random sampler during training to ensure balanced class representation.
Evaluation Metrics. We evaluate the models using the macro F1 score to account for the multi-class nature of the task, ensuring that performance is balanced across all bias categories. The final evaluation is conducted on the held-out test set. Additionally, we report the loss, runtime, and sample processing rates for completeness.
Results. After conducting a grid search across several models, we find that the roberta-base model achieves the best performance with an evaluation loss of 0.4035 and a macro F1 score of 0.9196 on the test set (see Table 1). Thus, roberta-base model is selected as the benchmark for political bias detection in subsequent experiments.
Table 1: Evaluation Results for Political Leaning Classifier
@@ 4.4 Generation Quality Metric
We introduced a metric for evaluating generation quality, specifically addressing the issue of repetitive content in later model iterations, which can skew traditional perplexity metrics. This metric is based on the Gibberish Detector (Jindal, 2021), which identifies incoherent or nonsensical text. The detector classifies text into four categories: (1) Noise, where individual words hold no meaning; (2) Word Salad, where phrases are incoherent; (3) Mild Gibberish, where grammatical or syntactical errors distort meaning; and (4) Clean, representing coherent, meaningful sentences. To evaluate generation quality, we assign the Gibberish score to each sentence: 3 for Clean, 2 for Mild Gibberish, 1 for Word Salad, and 0 for Noise. An average score is calculated for each article, resulting in the text quality index. It focuses on the coherence and meaning of the generated text and outperforms perplexity in accurately reflecting actual generation quality, as demonstrated in Section 5.2.
@@ 5 Results
In this section, we present the results on bias and generation performance of GPT-2 throughout iterative training, both with and without the deployment of mitigation strategies.
@@ 5.1 Political Bias
Figure 3: This figure displays the distribution of articles across the political bias labels—’Left’, ’Center’, and ’Right’—for GPT-2’s outputs, based on classifications made by the classifier (see Section 4.3).
Using an unbiased dataset of 1,518 human-written articles evenly distributed across three political-leaning categories, we generated 1,518 synthetic articles by deterministically predicting the next 64 tokens based on the previous 64 tokens. If GPT-2 had no pre-existing bias, the resulting distribution would be evenly spread across the three political-leaning labels. However, as shown in Figure 3, the model generates more center (47.9%) and right-leaning articles (46.8%), indicating a bias toward these categories prior to any fine-tuning.
Figure 4: This figure shows the percentage of right-leaning articles across multiple generations for three generation setups, with ’Synthetic’ representing the deterministic setup.
We then fine-tuned GPT-2 to produce Generation 0 and continued the process iteratively using synthetic datasets, up to Generation 10. As shown in Figure 4, the fine-tuned model displayed a significant amplification of bias toward the Right across all three generation setups after Generation 2. Interestingly, Generation 0 in the deterministic setup (the ’Synthetic’ line), which was fine-tuned on an unbiased dataset of real articles, produces an even higher percentage of right-leaning articles (54.5%) compared to the unfine-tuned GPT-2.
To ensure bias classification aligns with human intuition, we further examine model responses with human assistance. For illustration, we analyze a political opinion article, titled "First Read: Why It’s So Hard for Trump to Retreat on Immigration", from NBC News, a left-leaning outlet rated by AllSides (NBC News, 2016; AllSides, 2024b), critiquing President Trump’s immigration policies. Using a qualitative analysis grounded in foundational media studies (Entman, 1993; Rodrigo-Ginés et al., 2024; Groeling, 2013), we observe a significant rightward shift between generations 0 and 4 for this example, consistent with the classifier’s results (see Appendix C). As the generations progress, the synthetic texts increasingly frame Trump’s immigration policies as strong and effective. While the original article focuses on the dilemma and electoral interests behind Trump’s stance, Generation 0 begins to emphasize his determination and reliability, omitting the criticism present in the original. By Generation 4, the narrative shifts further, focusing almost entirely on portraying Trump’s personal qualities and electoral legitimacy, with statements like "he is not a politician, he is a man of action." Notably, from Generation 0 onward, the term "undocumented immigrant" in the original article is replaced with "illegal immigrants."
@@ 5.2 Generation Quality
Figure 5: This figure shows the average text quality index of synthetic articles, along with 95% confidence intervals, across generations for three generation setups, with ’Synthetic’ representing the deterministic setup.
As shown in Figure 5, there is a model collapse in terms of the deterioration of the generation quality across all three generation setups: the average text quality index (defined in Section 4.4) decreases over the iterative training. Moreover, the distributions of the text quality index for all three generation setups shift markedly toward the low-quality region and starts to produce data that is never produced by Generation 0, as shown by the distribution plots in Appendix D. These findings are consistent with the existing literature on model collapse, such as (Shumailov et al., 2024). However, we did not observe significant variation in variance, as seen in Figure 5 and distribution plots in Appendix D. In contrast, perplexity measurements show a steady decrease across generations for the deterministic and beam search generation setups, generally indicating improved generation quality (see Figure 11 in Appendix E).
For a closer look, as shown by the examples in Appendix F, the generated articles progressively lose coherence and relevance with each successive generation, as repetition and fragmented sentences become more common. By Generation 10, the text is largely incoherent and disconnected from the original content, diminishing its readability and meaning. However, despite the clear deterioration in generation quality, perplexity decreases over generations, as shown by the results at the end of each synthetic output example. This pattern is observed in most other synthetic outputs as well, suggesting that perplexity fails to accurately reflect the model’s true generation capabilities and may be biased toward repetitive content.
@@ 5.3 Mitigation Strategies
Figure 6: This figure shows the percentage of right-leaning articles across multiple generations for three mitigation strategies, with ’Synthetic’ representing the baseline deterministic setup.
Figure 7: This figure shows the average text quality index of synthetic articles, along with 95% confidence intervals, across multiple generations for three mitigation strategies, with ’Synthetic’ representing the baseline deterministic setup.
We applied three potential mitigation strategies under the deterministic generation setup: (1) Overfitting, which involved increasing the training epochs to 25 (five times the baseline) and setting weight decay to 0 to reduce regularization and encourage overfitting, as proposed by Taori and Hashimoto (2022) based on their uniformly faithful theorem of bias amplification; (2) Preserving 10% of randomly selected real articles during each round of synthetic fine-tuning, a method proposed and used in (Shumailov et al., 2024; Alemohammad et al., 2023; Dohmatob et al., 2024b; Guo et al., 2024); and (3) Accumulating all previous fine-tuning datasets along with the new synthetic dataset in each fine-tuning cycle, which was introduced by Gerstgrasser et al. (2024). As shown in Figure 6, overfitting helps reduce bias amplification in the early generations compared to the no-mitigation setup (the ’Synthetic’ line), but it fails to prevent bias amplification in the later generations. Additionally, it incurs a significant cost—further deterioration in generation quality, as shown in Figure 7. Notably, both the preservation and accumulation strategies perform well in mitigating both bias amplification and model collapse. Interestingly, these strategies even result in bias reduction.
@@ 5.4 Mechanistic Interpretation
To gain a clearer understanding of the causes of bias amplification and how it empirically differs from model collapse, we investigate how neurons behave and vary across different versions  [^2] We have 11 generations for each training setup and a total of 6 setups, resulting in 66 versions of fine-tuned GPT-2. of fine-tuned GPT-2 with differing levels of generation quality and bias performance. First, we examined how the weight of each neuron changes across different versions. For each of the 9,216 neurons, we calculated the correlation between its weight and the model’s bias performance (i.e. the percentage of right-leaning articles) across the 66 versions, as shown in Figure 12 in Appendix G. To statistically test the significance of the correlations, we calculated Newey-West adjusted p-values for the 9,216 neurons, using the null hypothesis that no significant relationship exists between the change in neuron weight and the change in bias performance, and applying the Bonferroni correction to maintain an overall 0.05 significance level. This yielded a set of 553 neurons whose weight value changes are significantly correlated with changes in bias performance (i.e. $p\text{-value}<1.36\times 10^{-6}$), indicating they are likely primary contributors to the shift in bias  [^3] All test results mentioned in this section are uploaded. In contrast, we identified only one neuron whose weight changes are significantly correlated with changes in generation quality. This suggests that shifts in bias performance and generation quality are likely driven by different underlying mechanisms. The formula for the statistical tests is provided in Appendix K.
Secondly, we examined how the activation values of each neuron change across different versions of GPT-2 when provided with the same input used to generate synthetic data (see Section 4.2). Specifically, we attached activation hooks to capture activation values (i.e., outputs from these layers during a forward pass) for 9,216 neurons. With the extracted activation values, we identified two sets: one consisting of 3,062 neurons whose activation value changes are significantly correlated with bias performance changes (i.e. $p\text{-value}<1.36\times 10^{-6}$), and another consisting of only 2 neurons whose activation value changes are significantly correlated with changes in generation quality.
6 Discussion
We now discuss the implications of bias amplification for the future development of large language models. As demonstrated, bias amplification is a parallel concept to model collapse and warrants equal emphasis in future research. In our case, we successfully identified ad hoc mitigation strategies, such as Preservation and Accumulation, which act as natural constraints on gradient updates during training by ensuring the model still considers the initial unbiased dataset. However, if the training dataset itself is biased or overrepresents certain demographic or political groups, these methods may fail, while pre-processing strategies could be too resource-intensive. Therefore, there is an urgent need to develop in-process mitigation strategies to address the constraint deficiency and promote more equitable and fair model development.
@@ 7 Limitations
While this work introduces a comprehensive framework for understanding bias amplification in large language models and provides empirical evidence using GPT-2, several limitations must be acknowledged. First, the scope of our experiments is restricted to political bias in the context of U.S. media. Additionally, our experiments were conducted using GPT-2, a relatively smaller model compared to state-of-the-art architectures like GPT-4 or LLaMA-2. Future research should extend our empirical approach to other contexts and larger LLMs.
Another limitation lies in our choice of mitigation strategies. While Preservation and Accumulation show promise in reducing bias amplification, their computational cost and scalability must be considered. Moreover, the mitigation strategies were tested primarily in the context of synthetic data generation, and their efficacy in real-world deployments requires further investigation.

### v2

@@ 4.1 Dataset Preparation
We randomly selected 1,518 articles from the Webis-Bias-Flipper-18 dataset (Chen et al., 2018), which contains political articles from a range of U.S. media outlets published between 2012 and 2018, along with bias ratings assigned at the time for each media source. These bias ratings, provided by AllSides, were determined through a multi-stage process incorporating assessments from both bipartisan experts and the general public (AllSides, 2024a). The random sampling was stratified based on bias ratings to ensure an even distribution of the 1,518 articles into three groups of 506 each, representing left-leaning, right-leaning, and center-leaning media.
4.2 Successive Fine-tuning
@@
Building on the approach in Shumailov et al. (2024); Dohmatob et al. (2024b), each training cycle begins with fine-tuning GPT-2 on a dataset of 1,518 news articles. The fine-tuned model, referred to as Generation 0, generates a synthetic dataset $D_{0}$ of the same size, which is then used for fine-tuning to produce Generation 1. This iterative process continues until Generation 10 is reached. Note that synthetic fine-tuning starts at Generation 1.
4.3 Synthetic Data Generation
Synthetic datasets, $\{D_{i}\}_{i=0}^{10}$, are generated as follows: Each original news article is tokenized into 64-token blocks, and for each block, the model predicts the next 64 tokens using deterministic generation to enhance the replicability of the study. The generated tokens are then decoded back into text, producing a synthetic dataset of the same size as the original.
@@ 4.4 Political Bias Metric
We develop a classification model to assess the political leaning of each LLM based on its generated synthetic news articles. The model is trained on the Webis-Bias-Flipper-18 dataset, excluding the 1,518 articles used for GPT-2 fine-tuning. To mitigate class imbalance, center-leaning articles are resampled to ensure equal representation across categories. The dataset is then divided into training (70%), validation (15%), and test (15%) subsets, stratified by bias label. Additionally, we conduct a human review to remove any identifiable information about media sources and authors.
After performing a grid search across multiple models, we find that roberta-base achieves the best performance, with an evaluation loss of 0.4035 and a macro F1 score of 0.9196 on the test set (see Table 1). Thus, we select roberta-base as the benchmark for political bias detection in subsequent experiments. Details on model training are provided in Appendix C.
Table 1: Evaluation Results for Political Bias Classifier
@@ 4.5 Generation Quality Metric
We introduce a metric to evaluate generation quality, specifically addressing the issue of repetitive content in later model iterations, which can distort traditional perplexity metrics. This metric is based on the Gibberish Detector (Jindal, 2021), which identifies incoherent or nonsensical text. The detector categorizes text into four levels: (1) Noise—individual words hold no meaning, (2) Word Salad—incoherent phrases, (3) Mild Gibberish—grammatical or syntactical distortions, and (4) Clean—coherent, meaningful sentences.
To quantify generation quality, each sentence receives a Gibberish score: 3 for Clean, 2 for Mild Gibberish, 1 for Word Salad, and 0 for Noise. The text quality index is computed as the average score across all sentences in an article. This metric prioritizes coherence and meaning, offering a better assessment of generation quality than perplexity.
@@ 5 Results
In this section, we analyze the evolution of political bias and generation quality in GPT-2 over successive iterations of synthetic fine-tuning, comparing results with and without mitigation strategies.
@@ 5.1 Political Bias
Figure 3: This figure displays the distribution of articles across the political bias labels—’Left’, ’Center’, and ’Right’—for GPT-2’s outputs, based on classifications made by Political Bias Metric.
GPT-2 was used to generate the synthetic dataset. Since the original human-written dataset is unbiased—with an equal number of articles for each political-leaning category—the synthetic dataset should ideally mirror this balanced distribution if GPT-2 had no pre-existing bias. Figure 3 presents the distribution of synthetic articles generated by GPT-2 across political bias labels. The model predominantly produces center-leaning (47.9%) and right-leaning (46.8%) articles, suggesting a pre-existing bias for these categories before any fine-tuning. Starting from the initial GPT-2 model, we fine-tuned it iteratively, generating synthetic datasets to train successive models up to Generation 10. Figure 4(a) illustrates how bias amplifies across generations. Surprisingly, fine-tuning on the unbiased real dataset increases right-leaning bias, with 53.7% of articles classified as right-leaning in Generation 0. Furthermore, without mitigation strategies, successive rounds of synthetic fine-tuning lead to a continuous rise in right-leaning articles, peaking at Generation 6 (67.6%) before
(a) Bias performance
(b) Generation quality
Figure 4: Evolution of bias performance and generation quality across generations, with Generation 0 representing GPT-2 fine-tuned on the unbiased dataset. The baseline (’Synthetic’) is compared with three mitigation strategies, and the text quality index is shown with 95% confidence intervals.
stabilizing. Figures 7 and 8 in Appendix D show the percentage of center-leaning and left-leaning articles across generations. Notably, the proportion of of center-leaning articles remains stable at approximately 35% throughout synthetic fine-tuning.
To further illustrate, we analyze how a specific article, "First Read: Why It’s So Hard for Trump to Retreat on Immigration", evolves in the synthetic generations. This case study reveals a progressive rightward shift in framing and word choice, mirroring the classifier’s results (Appendix E). As generations progress, the synthetic texts increasingly depict Trump’s immigration policies as strong and effective. While the original article highlights the dilemmas and electoral considerations behind Trump’s stance, Generation 0 begins to emphasize his determination and reliability, omitting the critical perspectives present in the original. By Generation 4, the narrative shifts even further, focusing almost entirely on portraying Trump’s personal qualities and electoral legitimacy, with statements such as "he is not a politician, he is a man of action." Notably, starting from Generation 0, the term "undocumented immigrant" in the original article is consistently replaced with "illegal immigrants."
@@ 5.2 Generation Quality
Figure 4(b) illustrates the text quality index across generations. In the training loop without any mitigation strategy, model collapse occurs, as evidenced by the gradual decline in the average text quality index. Furthermore, the distribution of the text quality index shifts significantly toward the lower-quality region over generations, eventually generating data that was never produced by Generation 0 (Figure 9 in Appendix F). These results align with prior research on model collapse, such as (Shumailov et al., 2024), though we did not observe substantial variation in variance. Conversely, perplexity measurements exhibit a consistent decline across generations, generally suggesting an improvement in generation quality (Figure 10 in Appendix G).
For a closer look, the examples in Appendix H illustrate how generated articles gradually lose coherence and relevance across generations, with increasing occurrences of repetition and fragmented sentences. By Generation 10, the text becomes largely incoherent and detached from the original content, reducing its readability and meaning. However, despite the evident decline in generation quality, perplexity decreases over generations, as indicated by the results at the end of each synthetic output example. This pattern is consistent across most synthetic outputs, suggesting that perplexity does not accurately capture the model’s true generative capabilities and is prone to artificial inflation in the presence of frequent repetitions.
@@ 5.3 Mitigation Strategies
We applied three mitigation strategies: (1) Overfitting, which involved increasing the training epochs to 25 (five times the baseline) and setting weight decay to 0 to reduce regularization and encourage overfitting, as proposed by Taori and Hashimoto (2022) based on the uniformly faithful theorem of bias amplification; (2) Preserving 10% of randomly selected real articles during each round of synthetic fine-tuning, a method proposed and used in
(a) Neuron weights vs. bias performance.
(b) Neuron weights vs. generation quality.
Figure 5: Scatter plots showing the Pearson correlation between neuron weights and model behavior.
(Shumailov et al., 2024; Alemohammad et al., 2023; Dohmatob et al., 2024b; Guo et al., 2024); and (3) Accumulating all previous fine-tuning datasets along with the new synthetic dataset in each fine-tuning cycle, which was introduced by Gerstgrasser et al. (2024). As shown in Figure 4(a), overfitting helps reduce bias amplification in the early generations compared to the no-mitigation setup (the ’Synthetic’ line), but it fails to prevent bias amplification in the later generations. Additionally, it incurs a significant cost—further deterioration in generation quality, as shown in Figure 4(b). Notably, both the preservation and accumulation strategies effectively mitigate model collapse and reduce bias, yielding 41.89% and 42.7% right-leaning articles, respectively, at Generation 10.
@@ 5.4 Mechanistic Interpretation
To gain a clearer understanding of the causes of bias amplification and how it empirically differs from model collapse, we investigate mechanistically how neurons behave and vary across different generations  [^3] We have 11 generations for each training setup and a total of 6 setups, resulting in 66 versions of fine-tuned GPT-2. of fine-tuned GPT-2 with differing levels of generation quality and bias performance. First, we examined how the weight of each neuron changes across different versions. For each of the 9,216 neurons, we calculated the correlation between its weight and the model’s bias performance (or generation quality) across the 66 versions, as shown in Figure 5.
To statistically test the significance of the correlations, we estimate the linear model:
$\Delta y_{i}=\alpha_{j}+\beta_{j}\Delta x_{i,j}+\epsilon_{i,j}$ (2)
where $\Delta y_{i}$ represents the change in the proportion of right-leaning articles (or text quality index) generated between model $i-1$ and model $i$, and $\Delta x_{i,j}$ denotes the corresponding change in the weight of neuron $j$ during the same transition. The coefficient $\beta_{j}$ captures the degree to which changes in the weight of neuron $j$ influence shifts in political bias (or generation quality), while $\alpha_{j}$ serves as a constant term, and $\epsilon_{i,j}$ represents the residual error. By applying a first-order difference to both $x_{i,j}$ and $y_{i}$, we mitigate potential serial correlations, ensuring that our regression estimates more accurately reflect the dynamic impact of individual neuron weight updates on bias amplification.
[…]
5.5 Further Investigation
(a) Bias performance
(b) Generation Quality
Figure 6: Evolution of bias performance and generation quality across generations, with Generation 0 representing GPT-2 fine-tuned exclusively on center-leaning news articles. The baseline (’Synthetic’) is compared with three mitigation strategies, and the text quality index is shown with 95% confidence intervals.
We conduct an alternative synthetic training cycle, beginning with GPT-2 fine-tuned on 1,518 randomly sampled center-labeled articles. We compare the baseline setup with the most effective and cost-efficient mitigation strategy identified in our previous results: Preservation. As shown in Figure 6, Preservation successfully prevents model collapse but fails to mitigate bias amplification in center-leaning article generation, which increases from 72.9% at Generation 0 to 88.2% at Generation 10. This result suggests that while reducing sampling error through techniques like Preservation effectively mitigates model collapse, it does not necessarily prevent bias amplification, thereby validating our mechanistic interpretation.
6 Discussion and Conclusion
We now explore the implications of our findings for the future research on large language models. Our results demonstrate that bias amplification operates through a distinct underlying mechanism from model collapse, as supported by both theoretical intuition and empirical evidence. Theoretically, the primary driven factor of model collapse is sampling error, but sampling error is only a auxiliary factor in bias amplification which means it is not a necessary or sufficient condition. Therefore, the mitigation strategy targeting sampling error is not necessarily helping with mitigating bias amplification. Empirically, we found mitigation strategies like preservation very effective in mitigating model collapse but failed at bias amplification in some cases. Even in cases that it helps with both, we do identify a distinct set of neurons responsible for the two phenomenons. Intuitively, the main reason for them to work on model collapse is, the preservation and accumulation propose a natural constraint on the learning process by recalling the real dataset in further synthetic training. However, when the real dataset itself is biased, the recalling behavior only raise up the dominance of biased patterns in the further training dataset. Indeed, applying bias-category-weighted sampling in preservation or accumulation strategies may help mitigate bias amplification. However, this approach inherently introduces additional sampling error, which could, in turn, incur model collapse. Thereby, This highlights the urgent need for more targeted and efficient mitigation strategies specifically addressing bias amplification to ensure fairer and more equitable model development.
To develop such targeted mitigation strategies, a deeper mechanistic understanding of bias amplification is essential. In our analysis, we adopt a statistical approach rather than Sparse Autoencoder (SAE) methods due to our focus on tracking the temporal dynamics of bias amplification across generations. This approach allows us to examine how neuron weights evolve over iterations and how these changes correlate with model bias, whereas existing SAE pipelines are primarily suited for static analysis. Additionally, political bias is a more nuanced concept compared to harmful or discriminatory outputs. It is characterized by the disproportionate overrepresentation or beautification of a particular political leaning’s ideas in a model’s generation. If content from different political perspectives is generated in a balanced manner, the model is not politically biased. Given this definition, pinpointing neurons responsible for such disproportionality using SAE is particularly challenging. Future research could focus on refining mechanistic analysis techniques for political bias and uncovering more effective ways to constrain bias amplification during synthetic model training.
@@ 7 Limitations
While this work introduces a comprehensive framework for understanding bias amplification in large language models and provides empirical evidence using GPT-2, several limitations must be acknowledged. First, the scope of our experiments is restricted to political bias in the context of U.S. media. The political spectrum may shift over time, necessitating periodic updates to the political bias classifier to ensure its accuracy when benchmarking recent datasets. Additionally, due to resource constraints, our experiments were conducted using GPT-2, a relatively small language model. Future work may extend our methodology to models with larger architectures. However, it’s important to note that employing a larger model, such as DeepSeek-V3, which relies on a considerably larger pre-training dataset and has many more parameters, will likely require larger synthetic fine-tuning datasets to effectively demonstrate its impact. For instance, given that our current fine-tuning datasets average 777,216 tokens, transitioning from GPT-2 to DeepSeek-V3 could necessitate much larger datasets and incur significant fine-tuning costs.
Another limitation lies in our choice of mitigation strategies. While Preservation and Accumulation show promise in reducing model collapse, their computational cost and scalability must be considered. Moreover, the mitigation strategies were tested primarily in the context of synthetic fine-tuning, and their efficacy in real-world deployments requires further investigation.